OpenStudy (anonymous):

Medals!!!! :D Determine two pairs of polar coordinates for the point (2, -2) with 0° ≤ θ < 360°.

OpenStudy (anonymous):

Answer choices: a. (2 square root of 2, 225°), (-2 square root of 2, 45°) b. (2 square root of 2, 135°), (-2 square root of 2, 315°) c.(2 square root of 2, 315°), (-2 square root of 2, 135°) d.(2 square root of 2, 45°), (-2 square root of 2, 225°)

OpenStudy (anonymous):

@TheStarlingHasFlown @amistre64 ^.^

OpenStudy (anonymous):

@Michele_Laino @Keygrover @BloomLocke367 somebody help me please....

OpenStudy (michele_laino):

we have to apply these formulas, in order to change from cartesian coordinates (x,y) to polar coordinates (r, \theta): \[\left\{ \begin{gathered} x = r\cos \theta \hfill \\ y = r\sin \theta \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):

from those formulas, we get: \[r = \sqrt {{x^2} + {y^2}} \]

OpenStudy (anonymous):

okay....

OpenStudy (michele_laino):

in our case, we have: \[r = \sqrt {{x^2} + {y^2}} = \sqrt {{{\left( { - 2} \right)}^2} + {2^2}} = ...?\]

OpenStudy (michele_laino):

what is r?

OpenStudy (michele_laino):

\[r = \sqrt {{x^2} + {y^2}} = \sqrt {{2^2} + {{\left( { - 2} \right)}^2}} = \sqrt {4 + 4} = ...?\]

OpenStudy (anonymous):

2sqrt2??

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

oh :)

OpenStudy (michele_laino):

now, starting from my transformation formulas, we get: \[\Large \tan \theta = \frac{y}{x}\]

OpenStudy (michele_laino):

so, in our case , we have: \[\Large \tan \theta = \frac{y}{x} = \frac{{ - 2}}{2} = ...?\]

OpenStudy (michele_laino):

what is the value of \[\tan \theta \]

OpenStudy (anonymous):

\[\tan \theta-1??\]

OpenStudy (michele_laino):

right! \[\tan \theta = \frac{y}{x} = \frac{{ - 2}}{2} = - 1\]

OpenStudy (anonymous):

oh! :) i'm not used to being right in math!! X)

OpenStudy (anonymous):

i wasn't sure if i understood exactly what you were asking

OpenStudy (michele_laino):

so, what is theta? please keep in mind that drawing: